Week of 190311

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local: Miro (Chair), Joao (Storage), Gavin (Compute), Maarten (WLCG), Alberto (Monitoring), Belinda (Storage)
  • remote: John (RAL),. Stephan (CMS), Alexander (NL-T1), Di (TRIUMF), David (IN2P3), Carlos (ATLAS), Elena (CNAF), Xin (BNL), Sang Un (KISTI), Federico (LHCb), Dave (FNAL), Jose (PIC)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Smooth running at 300k job slots
    • ATLAS EOS downtime tomorrow. It would need an outage entry in GOCDB.
    • Check for FTS monitoring, is it using the correct ATLAS VO feed?

  • CMS reports ( raw view) -
    • No major grid computing issue(s)
    • Tier-2 US xrootd authentication failure for certificates proxied Sunday early morning UTC (day-light-savings bug?)
    • Running smoothly at about 220k cores
      • about 50k cores analysis, the rest production

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity
      • User jobs and MC productions
    • Issues
      • CERN: Some tickets for CERN/EOS are still open, even though problems mostly gone. Not clear why.

Sites / Services round table:

  • ASGC: NC
  • BNL: NTR
  • EGI: NC
  • IN2P3: Schedule maintenance on March 19th and 20th due to work on air conditioning.
    • CEs in downtime for 2 days
    • SEs in downtime on March 19th only
  • JINR: Temporary loss of 84 10-core slots due to h/w problem, vendor is investigating the problem.
  • KISTI: Site reported issue with IPv6 storage.
  • KIT: NTR
  • NDGF: Tuesday is for dCache pool upgrades at NDGF. Shouldn't cause big issues. Some data might become shortly unavailable.
  • NL-T1: NTR
  • NRC-KI: NC
  • OSG: NC
  • PIC: NTR
  • RAL: NTR

  • CERN computing services: NTR
  • CERN storage services:
    • CERN AFS external disconnection test for one week on 3 April 2019 from 9:00. More info here: OTG:0048585
  • CERN databases: NTR
  • Monitoring: NTR
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR


  • NOTE: next week the Joint HSF / OSG / WLCG Workshop will be held at JLab.
  • The operations meeting next Mon will be virtual .
  • You may provide relevant incidents, announcements etc. for the operations record.
Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2019-03-12 - MaartenLitmaath
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback